# A General Class of Model-Free Dense Precision Matrix Estimators

### v1

@@ 4.3 Consistent Estimation of Rows
We provide our theorem for estimating rows of the precision matrix. This shows that each row of the precision matrix in general case can be estimated, uniformly over all rows, consistently through our estimator. Specific results regarding adaptive PCR and RRE will be discussed in subsections 4.3.1-4.3.2. Note that the main diagonal terms rate determine the rate of convergence. The off-diagonal terms estimation is consistent but converges at a faster rate than the main diagonal terms. This can be seen through comparing the rates in Theorem 1 and via proof of Theorem 3 and Remark-Item 4 in Theorem 2. Denote each row of $\Theta$ as $\Theta_{j}^{\prime}$, which is a $1\times p$ vector. As an example, first row can be represented as $\Theta_{1}^{\prime}:=(\frac{1}{\tau_{1}^{2}},\frac{-\alpha_{1}^{*^{\prime}}}{\tau_{1}^{2}})$. Let $\Theta_{j}$ be the column representation of the row $\Theta_{j}^{\prime}$. Let the row estimator be $\tilde{\Theta}_{j}^{\prime}$, and for the first row this is $\tilde{\Theta}_{1}^{{}^{\prime}}:=(\frac{1}{\tilde{\tau}^{2}},\frac{-\tilde{\alpha}_{1}^{{}^{\prime}}}{\tilde{\tau}_{1}^{2}})$.
Theorem 3.
Under Assumptions 1-5, with $r_{n}/\delta_{n}\to 1$
$\max_{1\leq j\leq p}\|\tilde{\Theta}_{j}^{\prime}-\Theta_{j}^{\prime}\|_{2}$ $=$ $O_{p}(K\frac{\sqrt{plogp}}{\sqrt{n}}r_{w1})+O_{p}(d_{2n}r_{w2})$
@@
where $r_{w1},r_{w2}$ are defined in (9)(10).
Remark 7.
@@
  • We see that through definition of $r_{w1},r_{w2}$ one of the keys to consistent estimation is the ratio $K/\bar{\xi}$, and this ratio needs to go to zero. So we cannot have a large hidden factor system and expect consistency. Key is the tradeoff between two terms. With low signal-to-noise ratio, the number of hidden factors has to be small.
  • Note that the rate for Theorem 3 is obtained for all linear estimators in Section 3, except RRE when $p-1>n$, the interpolating estimator. This is explained in detail, below in Corollary 6 and Appendix Corollary OA.3.
@@ 4.3.1 Adaptive PCR Estimator
In case of adaptive PCR, the random terms in the upper bound in Theorem 3, can be bounded as: $\bar{r}\leq K$ wpa1, $\bar{\Psi}/\bar{\eta}\leq 1$ wpa1, $\bar{\eta}\geq\delta_{n}$ with wpa1. So Theorem 3 rate can be simplified in adaptive PCR, with $\hat{\Theta}_{PCR,j}^{\prime}$ representing the estimator of $j$ th row of $\Theta$. The following Corollary follows from Theorem 3 by replacing $r_{w1},r_{w2}$ with $r_{w1,PCR},r_{w2,PCR}$ rates that are shown before Corollary OA.1 in Online Appendix.
Corollary 4.
Let $p/n\to[0,\infty)$ and with Assumptions 1-5, and with $r_{n}/\delta_{n}\to 1$
$\max_{1\leq j\leq p}\|\hat{\Theta}_{PCR,j}^{\prime}-\Theta_{j}^{\prime}\|_{2}$ $=$ $O_{p}\left(K\sqrt{logp}\max(\frac{\sqrt{logn+K}}{\sqrt{n\delta_{n}}},\frac{\sqrt{K}}{\sqrt{\bar{\xi}\delta_{n}}})\right)$
$+$ $O_{p}\left(d_{2n}\max(\frac{\sqrt{logn+K}}{\sqrt{n\delta_{n}}},\frac{\sqrt{K}}{\sqrt{\bar{\xi}}})\right)$
@@
  • We analyze how consistency can be achieved in this adaptive PCR. We start with Assumption 5(iii) which may be the most stringent term for consistency. We use example (8) in Assumption 5(iii) explanation to illustrate our point. Take the case of $p/n\to\gamma$, with $\gamma\in(1,\infty)$, $\gamma$ is a positive constant. Case of $p/n\to\gamma$, with $\gamma\in[0,1]$ is similar hence skipped. What if we increase $n,p$ together but keep $\gamma$ fixed. Consistency will require,
    $\frac{K\sqrt{log(\gamma n)}\sqrt{logn+K}}{\sqrt{n\delta_{n}}}\to 0,$ (22)
    which is possible with Assumption 5(ii) imposing $K=o(n^{1/3})$, and since $\delta_{n}$ is nondecreasing in n . Then we need
    $\frac{K^{3/2}\sqrt{log(\gamma n)}}{\sqrt{\delta_{n}\bar{\xi}}}\to 0$ (23)
    which can be possible with (8), $r_{n}/\delta_{n}\to 1$ and since $\frac{K^{3/2}\sqrt{log(\gamma n)}}{d_{1n}}\to 0,$ with increasing $d_{1n}$ with $n$, and can grow larger than $K^{3/2}\sqrt{logn}$. Next, we consider in Assumption 5(iii) $d_{2n}\sqrt{\frac{logn+K}{n\delta_{n}}}\to 0,$ since $K=o(n^{1/3})$ and $d_{2n}$ can diverge with n but not too fast (i.e. such as a slowly varying. function in n). Then for the second part of that consistency condition we also need with example (8) $d_{2n}\sqrt{\frac{K}{\bar{\xi}\delta_{n}}}\leq\frac{d_{2n}K^{1/2}}{d_{1n}}\to 0.$ So the signal $d_{1n}$ can be large and bring consistency.
  • An interesting connection of the above theorem’s results and Fan et al. (2013) can be made as well. First, one recalls that Fan et al. (2013) estimates hidden factors through principal components and feeds such estimates into covariance matrix of returns which is then inverted via Sherman-Woodbury- Morrison formula. However, underlying hidden factor rotation results in having favorable rates of convergence provided that p is large.
@@ 4.3.2 Ridgeless Estimators
When we consider the sub case of RRE with $p<n-1$, we set wpa1 $\bar{r}<p$, $\bar{\Psi}=0$, and $\bar{\eta}\geq c$, where $c$ is a positive constant in Theorem 3 result above. Let the $j$ th row of estimator of the precision matrix be $\hat{\Theta}_{RRE,j}^{\prime}$. To obtain Corollary 5, we use Corollary OA.2 and Corollary 2 with $r_{w1},r_{w2}$ in Theorem 3 replaced with $r_{RRE-l}$ rate after Corollary OA.1 in Online Appendix.
Corollary 5.
Set $p/n=a_{2n}\to 0$, and under Assumptions 1-5, and $r_{n}/\delta_{n}\to 1$
@@
Remark 9.
  • Note that unless $p/n\to 0$, it is not clear that RRE will be consistent. This is a major handicap compared with adaptive PCR, since adaptive PCR is consistent even when $p/n=\gamma\in(0,1)$. Also we see that with $K<p$, the rate of convergence of adaptive PCR in case of $p/n\to 0$ will be better than RRE here. Only in finite samples RRE, may have an advantage due to nuisance parameter $\bar{\Psi}=0$ in RRE. RRE can work well only when $p$ is very small compared with $n$. Also, when $p-1<n$ there is a notable advantage of RRE. Namely, it can be run without tuning parameter as RRE is equivalent to OLS estimator in this case.
  • One remaining question is with $p/n=a_{2n}\to 0$ then which method will have a better rate. Consider the first part of Assumption 5(iii), then in RRE
    $a_{2n}^{1/2}K(\sqrt{logp})r_{RRE-l}=a_{2n}^{1/2}K\sqrt{logp}\max(\sqrt{\frac{logn+p}{n}},\sqrt{\frac{K}{\bar{\xi}}}),$
    and in adaptive PCR the same rate is
    $a_{2n}^{1/2}K(\sqrt{logp})r_{w1,PCR}=a_{2n}^{1/2}K\sqrt{logp}\max(\sqrt{\frac{logn+K}{n\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}\delta_{n}}}),$
    which will be smaller since $K<p$ and $\delta_{n}$ is nondecreasing in n. The second part of Assumption 5(iii) follows a similar pattern and hence adaptive PCR is a better estimator asymptotically than RRE when $p-1<n$.
Next, we consider RRE estimator when $p-1>n$. Note that in this scenario main diagonal term estimation is done differently since RRE is an interpolating estimator. We prove this after Corollary OA.2. In the proof of Theorem 3, it is clear that the rate of convergence should be the slower of the main diagonal or off-diagonal terms. Let the j-th row of the precision matrix be $\hat{\Theta}_{INT,j}$. Corollary 6 below proves $\hat{\Theta}_{INT,j}$’s consistency by combining Corollaries 3 and OA.3.
Corollary 6.
@@
$+$ $O_{p}\left(\max(\frac{K^{3/2}}{n^{1/2}},K\frac{\sqrt{logp)}}{n^{1/2}})\right)=o_{p}(1).$
Remark 10.
  • First of all we still get consistency with $p-1>n$, as long as signal-to-noise ratio dominates number of factors multiplied by $d_{2n}$. This can be seen easily by $K\sqrt{logp/n}\to 0$, and also we need signal-to-noise condition: $\sqrt{K}d_{2n}/\sqrt{\bar{\xi}}\to 0$.
@@
  • As described in Remark 5-item 2 of Corollary 3 and item 1 above, double descent is possible in RRE in precision matrix estimation. This extends the concept of double descent from linear regression to precision matrix estimation. This may have important consequences in empirics. We show one of these findings in Sharpe Ratio in US Stock market.
4.4 Uses of Precision Matrix Estimate
Here in this section we discuss the usage of our precision matrix estimator in three different cases of interest in economics and finance. In addition to the examples that we will discuss below, we should also add that many of our results are of independent statistical interest as well. We single out two particular features of our statistical analyses, (i) lack of exact sparsity assumption between outcomes with model-free dense estimators and (ii) an explicit connection of the estimation error to signal-to-noise ratio, as results that significantly advance the current state of the art in the relevant statistics precision matrix estimation related literature.
In financial econometrics and finance, one of the key issues is large portfolio formation. Define the number of assets of a large portfolio as $p$, and the sample size as $n$. Two precision matrix estimating approaches distinguished themselves in recent literature: (i) the first one that relies on the estimation and inversion of the covariance matrix (see, e.g., Fan et al. (2013) and Ledoit and Wolf (2017) where covariance matrix is obtained via hidden factor model and shrinkage estimation, respectively); and (ii) the second one that directly estimates the precision matrix and relies on the so-called nodewise regression approach (see, e.g., Callot et al. (2021) and Caner et al. (2023) for nodewise regression with observed factors). We introduce a new estimator that can be used to consistently estimate the Sharpe Ratio even when $p>n$ (Sharpe Ratio is one of the key practical financial metrics as it can measure return-risk ratio for a given portfolio).
The second possibility is a simple modification of our method can be used to estimate inverted Gram matrix thereby replacing the so-called relaxed inverse typically used to get a high dimensional linear regression parameters confidence intervals (see, e.g., van de Geer et al. (2014)). Our idea can also be used to form and change the debiased high dimensional estimators used in van de Geer et al. (2014) and Caner and Kock (2018). Below, we provide some details on how this can be achieved.
The third possibility for the utilization of our methodology is the network analysis. One first recalls that nodewise regression can be tied to network analysis (see, e.g., Meinshausen and Bühlmann (2006). Given the connection between our estimator and nodewise regression, one then observes that our estimator fits the same network structure but with an advantage of not needing to impose sparsity on the links between two nodes.
We have three examples of our estimators, and the first one is Sharpe-Ratio analysis, and the second one is the de-biased high dimensional estimator formation for the confidence intervals for structural parameters in cross section analysis, and the third one is network analysis. In this part, first, we provide a Sharpe Ratio metric that use precision matrix of asset returns. The metric is the Sharpe ratio (squared) from the maximum Sharpe Ratio portfolio. There are other metrics as discussed heavily in Caner et al. (2023) for example, but the emphasis of our paper is introduction of a new class of estimators for the precision matrix not the detailed analysis of financial metrics. We also think that the analysis of financial metrics can be done in another paper with this approach, since the proofs will be non-standard and will take a larger space with very detailed analysis through several matrix norm inequalities. The metric below is chosen since the contribution of the precision matrix is very clear to the financial metric. To obtain the (squared) Sharpe ratio from the maximum Sharpe Ratio portfolio we maximize
$w_{MS}^{*}:=argmax_{w\in R^{p}}w^{\prime}\mu\quad{\mbox{s}ubject}\quad{\mbox{t}o}\quad w^{\prime}\Sigma w\leq\sigma^{2},$
where $\sigma^{2}$ is the variance of the portfolio. The optimal solution is $w_{MS}^{*}=\frac{\sigma\Sigma^{-1}\mu}{\sqrt{\mu^{\prime}\Sigma^{-1}\mu}}.$ Substituting this into Sharpe Ratio formula we have $SR_{MS}=\frac{w_{MS}^{*^{\prime}}\mu}{w_{MS}^{*^{\prime}}\Sigma w_{MS}^{*}}=\sqrt{\mu^{\prime}\Sigma^{-1}\mu}.$ The estimator is: $\widehat{SR}_{MS}^{2}:=\hat{\mu}^{\prime}\tilde{\Theta}^{\prime}\hat{\mu},$ with $\hat{\mu}:=\frac{1}{n}\sum_{t=1}^{n}Y_{t}$.
This can be found in Caner et al. (2023) as well as several other papers in the literature, and used in practice. The prevalent approaches in present day literature impose sparsity either on the covariance or on the precision matrix of errors, and prove consistency results relying on those assumptions. Without exact sparsity, the proofs will not be valid for $p>n$ case. On the other hand, our methodology allows to circumvent the sparsity assumption. The second issue is that rates of convergence are usually only tied to the number of assets, p, sample size, n, and the number of factors, K. Our methodology allows to additionally tie them to signal-to-noise ratio as well.
In our second example, we show how our precision matrix can be modified-used to form a debiased lasso estimator for estimating structural parameters, and this type of estimator can then be used to have confidence intervals for these structural parameters. Our setup is a simple linear high dimensional linear regression as in Caner and Kock (2018) and van de Geer et al. (2014).
$Z=Y\beta+\epsilon,$
where $Z$ represents the outcome variables here, and it is $n\times 1$ vector, where $Z_{i},i=1,\cdots,n$ are cross-section units (instead of $t$ index in the other parts of the paper, we use $i$ for cross section data), and $Y:n\times p$ matrix, where each row represents cross section units, and $j=1,\cdots,p$ represent regressors. Let $\beta:p\times 1$ be the structural parameters of interest, and let $s_{0}$ of them to be nonzero. We allow $s_{0}$ n to be nondecreasing in $n$. Let $\epsilon:n\times 1$ represent the errors. Also let $EY_{i}\epsilon_{i}=0$, and $\epsilon_{i}$ is iid as well as $Y_{i}$. For example, $Z$ can represent average (over time) growth rate of GDP of a $n$ group of countries, and each column of $Y$ represents each regressor, such as public debt to GDP ratio, current account deficit, inflation, that contributes to this average GDP growth rate. We also assume existence of the following relation
$Y=FA+U,$
as can be formed by (4)(5), so that regressors are governed by latent factors. Note that using lasso estimation of $\beta$ results in biased estimates and is not uniformly consistent, hence any confidence intervals will not be valid and not asymptotically honest as discussed in Caner and Kock (2018). To solve this issue, a method is debiasing the lasso estimators. However, debiased lasso estimators depend on the estimation of the inverse of the population inner-product $EY_{i}Y_{i}^{\prime}$, where $Y_{i}:p\times 1$ vector, (row of $Y$ represented in column form). It is not possible to use sample $Y^{\prime}Y/n$ and invert it to get an estimate since if $p>n$, $Y^{\prime}Y/n$ is singular. First we show how to form $[EY_{i}Y_{i}^{\prime}]^{-1}$ which is a $p\times p$ matrix. This can be formed exactly with the analysis in Appendix B, (with only imposing $\mu_{j}=0,\mu_{-j}=0_{p-1}$, where $0_{p-1}$ vector is a p-1 vector of zeroes in each element). Also see p.157 of Caner and Kock (2018) for this as well. Our estimator for $\tilde{\Theta}$ is exactly the same as in Section 4 with no change. Then the debiased estimator, $\hat{b}$, can be formed through, using the analysis in Section 3.1, equations (7)-(9) of Caner and Kock (2018) but with our new estimator of inverse of $EY_{i}Y_{i}^{\prime}$ rather than the nodewise regression in Caner and Kock (2018),
$\hat{b}=\hat{\beta}+\tilde{\Theta}Y^{\prime}(Z-Y\hat{\beta})/n,$
where $\hat{\beta}$ is the lasso estimator for $\beta$ by regressing a $l_{1}$ penalized regression of $Z$ on $Y$. We want to focus on this estimator and how to form confidence intervals for our next project.
Our network estimator can be based on the theory in Appendix B, and it can be used, a detailed analysis is beyond the scope of this paper. But a start is the social interaction matrix estimation in De Paula et al. (2024).
@@ 5 Empirical Application
The objective of the empirical application is to conduct an out-of-sample comparison among estimations. We estimate Maximum Sharpe Ratio (MSR) portfolio using estimators, and use four commonly used financial metrics: the Sharpe Ratio (SR), average return, portfolio turnover, and portfolio variance. We use data from January 1995 to December 2023, with an out-of-sample period from January 2015 to December 2023. We selected all stocks included in the S&\&&P 500 for at least one month during the out-of-sample period, ensuring the data were available for the entire 1995-2023 period, resulting in 357 stocks. The out-of-sample period includes stock market of 2022 when S&\&&P 500 index was down by 18.1 %.
@@ 5.1 Metrics
We use a rolling window set up for the out-of-sample estimation of the Sharpe Ratio following Callot et al. (2021) and Caner et al. (2023). The sample size n is divided into in-sample $(1:n_{I})$ and out-of-sample $(n_{I}+1:n)$. Initially, we estimate the portfolio $\hat{w}_{n_{I}}$ in the in-sample period and the out-of-sample portfolio returns $\hat{w}^{\prime}_{n_{I}}y_{n_{I}+1}$. Then, we roll the window forward by one element $(2:n_{I}+1)$, forming a new in-sample portfolio $\hat{w}_{n_{I}+1}$ and caculating out-of-sample portfolio returns $\hat{w}_{n_{I+1}}^{\prime}y_{n_{I}+2}$. This process is repeated until the end of the sample.
The out-of-sample average return and variance without transaction costs are computed as follows: $\hat{\mu}_{os}=\frac{1}{n-n_{I}}\sum_{t=n_{I}}^{n-1}\hat{w}_{t}^{\prime}y_{t+1},\hat{\Sigma}^{2}_{y,os}=\frac{1}{n-n_{I}-1}\sum_{t=n_{I}}^{n-1}(\hat{w}_{t}^{\prime}y_{t+1}-\hat{\mu}_{os})$. We estimate the Sharpe Ratios without transaction costs: $SR=\frac{\hat{\mu}_{os}}{\hat{\Sigma}_{y,os}}$
When transaction costs are present, the definition of the metrics is modified. Let $c$ be the transaction cost, which we set 10 basis points as. Let $y_{P,t+1}=\hat{w}_{t}^{\prime}y_{t+1}$ be the return of the portfolio $P$ in period $t+1$; in the presence of transaction costs, the return is defined as $y^{Net}_{P,t+1}=y_{P,t+1}-c(1+y_{P,t+1})\sum_{j=1}^{p}|\hat{w}_{j,t+1}-\hat{w}^{+}_{j,t}|$ where $\hat{w}^{+}_{j,t}=\frac{\hat{w}_{j,t}(1+y_{j,t+1})}{1+y_{P,t+1}}$, with $y_{j,t+1}$ representing the excess return added to risk-free-rate for $j$th, and $y_{P,t+1}$ representing the portfolio $P$ excess return added to the risk-free rate. The adjustment made in $\hat{w}^{+}_{j,t}$ is because the portfolio at the end of the period has changed compared to the portfolio at the beginning of the period. Then, we define the mean and variance of the portfolio returns with transaction costs as: $\hat{\mu}_{os-c}=\frac{1}{n-n_{I}}\sum_{t=n_{I}}^{n-1}y^{Net}_{P,t}\mbox{ and }\hat{\Sigma}^{2}_{y,os-c}=\frac{1}{n-n_{I}-1}\sum_{t=n_{I}}^{n-1}(y^{Net}_{P,t}-\hat{\mu}_{os-c})^{2}$. The portfolio returns are replaced by the returns with transaction costs when we calculated the Sharpe Ratio with transaction costs. Finally, portfolio turnover is defined as: $\text{turnover}=\sum_{j=1}^{p}|\hat{w}_{j,t+1}-\hat{w}_{j,t}^{+}|.$

### v2

@@ 4.3 Consistent Estimation of Rows
We provide our theorem for estimating rows of the precision matrix. This shows that each row of the precision matrix in general case can be estimated, uniformly over all rows, consistently through our estimator. Specific results regarding adaptive PCR and RRE will be discussed in subsections 4.3.1-4.3.2. Note that the main diagonal terms rate determine the rate of convergence. The off-diagonal terms estimation is consistent but converges at a faster rate than the main diagonal terms. This can be seen through comparing the rates in Theorem 1 and via proof of Theorem 3 and Remark 2 in Theorem 2. Denote each row of $\Theta$ as $\Theta_{j}^{\prime}$, which is a $1\times p$ vector. As an example, first row can be represented as $\Theta_{1}^{\prime}:=(\frac{1}{\tau_{1}^{2}},\frac{-\alpha_{1}^{*^{\prime}}}{\tau_{1}^{2}})$. Let $\Theta_{j}$ be the column representation of the row $\Theta_{j}^{\prime}$. Let the row estimator be $\tilde{\Theta}_{j}^{\prime}$, and for the first row this is $\tilde{\Theta}_{1}^{{}^{\prime}}:=(\frac{1}{\tilde{\tau}^{2}},\frac{-\tilde{\alpha}_{1}^{{}^{\prime}}}{\tilde{\tau}_{1}^{2}})$.
Theorem 3
Under Assumptions 1-5,
$\max_{1\leq j\leq p}\|\tilde{\Theta}_{j}^{\prime}-\Theta_{j}^{\prime}\|_{2}$ $=$ $O_{p}(K\frac{\sqrt{plogp}}{\sqrt{n}}r_{w1})+O_{p}(d_{2n}r_{w1})$
@@
where $r_{w1}$ is defined in (9).
Remark 3
@@
  • We see that through definition of $r_{w1}$ one of the keys to consistent estimation is the ratio $K/\bar{\xi}$, and this ratio needs to go to zero. So we cannot have a large hidden factor system and expect consistency.
  • Note that the rate for Theorem 3 is obtained for all linear estimators in Section 3, except RRE when p1>np-1>n, the interpolating estimator. This is explained in detail, below in Corollary 6 and Online Appendix Corollary A.3.
@@ 4.3.1 Adaptive PCR Estimator
In case of adaptive PCR, the random terms in the upper bound in Theorem 3, can be bounded as: $\bar{r}\leq K$ wpa1, $\bar{\Psi}/\bar{\eta}\leq 1$ wpa1, $\bar{\eta}\geq\beta$ with wpa1. So Theorem 3 rate can be simplified in adaptive PCR, with $\hat{\Theta}_{PCR,j}^{\prime}$ representing the estimator of $j$ th row of $\Theta$. The following Corollary follows from Theorem 3 by replacing $r_{w1}$ with $r_{w1,PCR}$ rate that is shown before Corollary A.1 in Online Appendix.
Corollary 4
Let $p/n\to[0,\infty)$ and with Assumptions 1-5,
$\max_{1\leq j\leq p}\|\hat{\Theta}_{PCR,j}^{\prime}-\Theta_{j}^{\prime}\|_{2}$ $=$ $O_{p}\left(K\sqrt{logp}\max(\frac{\sqrt{logn+K}}{\sqrt{n}},\frac{\sqrt{K}}{\sqrt{\bar{\xi}}})\right)$
$+$ $O_{p}\left(d_{2n}\max(\frac{\sqrt{logn+K}}{\sqrt{n}},\frac{\sqrt{K}}{\sqrt{\bar{\xi}}})\right)$
@@
  • We analyze how consistency can be achieved in this adaptive PCR. It can be shown that with $p>n$ and with both $p,n$ growing we need signal-to-noise ratio growing and $d_{2n}$ to grow slowly in $n$.
@@ 4.3.2 Ridgeless Estimators
When we consider the sub case of RRE with p<n1p<n-1, we set wpa1 $\bar{r}<p$, $\bar{\Psi}=0$, and $\bar{\eta}\geq c$, where $c$ is a positive constant in Theorem 3 result above. Let the $j$ th row of estimator of the precision matrix be $\hat{\Theta}_{RRE,j}^{\prime}$. To obtain Corollary 5, we use Corollary A.2 and Corollary 2 with $r_{w1}$ in Theorem 3 replaced with $r_{RRE-l}$ rate after Corollary A.1 in Online Appendix.
Corollary 5
Set $p/n=a_{2n}\to 0$, and under Assumptions 1-5,
@@
Remark 5
  • Note that unless $p/n\to 0$, RRE will not be consistent.
Next, we consider RRE estimator when p1>np-1>n. Note that in this scenario main diagonal term estimation is done differently since RRE is an interpolating estimator. We prove this after Corollary A.2. In the proof of Theorem 3, it is clear that the rate of convergence should be the slower of the main diagonal or off-diagonal terms. Let the j-th row of the precision matrix be $\hat{\Theta}_{INT,j}$. Corollary 6 below proves $\hat{\Theta}_{INT,j}$’s consistency by combining Corollaries 3 and A.3.
Corollary 6
@@
$+$ $O_{p}\left(\max(\frac{K^{3/2}}{n^{1/2}},K\frac{\sqrt{logp}}{n^{1/2}})\right)=o_{p}(1).$
Remark 6
  • First of all we still get consistency with p1>np-1>n, as long as signal-to-noise ratio dominates number of factors multiplied by $d_{2n}$. This can be seen easily by $K\sqrt{logp/n}\to 0$, and also we need signal-to-noise condition: $\sqrt{K}d_{2n}/\sqrt{\bar{\xi}}\to 0$.
@@
  • Double descent is possible in RRE in precision matrix estimation. Large $p,n$ can result in increasing size of the signal, and signal-to-noise ratio $\delta_{n},\bar{\xi}$ when $p>n$ respectively. Specifically when $\delta_{n}>logp$ we can observe double descent. With $p<n$, the consistency is achieved with $p/n\to 0$. This extends the concept of double descent from linear regression to precision matrix estimation. This may have important consequences in empirics. We show one of these findings in Sharpe Ratio in US Stock market.
@@ 5 Empirical Application
The objective of the empirical application is to conduct an out-of-sample comparison among estimations. We estimate Maximum Sharpe Ratio (MSR) portfolio using estimators, and use four commonly used financial metrics: the Sharpe Ratio (SR), average return, portfolio turnover, and portfolio variance. We use data from January 1995 to December 2023, with an out-of-sample period from January 2015 to December 2023. We selected all stocks included in the S&\&P 500 for at least one month during the out-of-sample period, ensuring the data were available for the entire 1995-2023 period, resulting in 357 stocks. The out-of-sample period includes stock market of 2022 when S&\&P 500 index was down by 18.1 %.
@@ 5.1 Metrics
We use a rolling window set up for the out-of-sample estimation of the Sharpe Ratio following Callot et al. (2021) and Caner et al. (2023). The sample size n is divided into in-sample $(1:n_{I})$ and out-of-sample $(n_{I}+1:n)$. Initially, we estimate the portfolio $\hat{w}_{n_{I}}$ in the in-sample period and the out-of-sample portfolio returns $\hat{w}^{\prime}_{n_{I}}y_{n_{I}+1}$. Then, we roll the window forward by one element $(2:n_{I}+1)$, forming a new in-sample portfolio $\hat{w}_{n_{I}+1}$ and calculating out-of-sample portfolio returns $\hat{w}_{n_{I+1}}^{\prime}y_{n_{I}+2}$. This process is repeated until the end of the sample.
The out-of-sample average return and variance without transaction costs are computed as follows: $\hat{\mu}_{os}=\frac{1}{n-n_{I}}\sum_{t=n_{I}}^{n-1}\hat{w}_{t}^{\prime}y_{t+1},\hat{\Sigma}^{2}_{y,os}=\frac{1}{n-n_{I}-1}\sum_{t=n_{I}}^{n-1}(\hat{w}_{t}^{\prime}y_{t+1}-\hat{\mu}_{os})^{2}$. We estimate the Sharpe Ratios without transaction costs: $SR=\frac{\hat{\mu}_{os}}{\hat{\Sigma}_{y,os}}$
[…]